# Physics motivated models of pulsar X-ray hotspots: off-center dipole configurations

### v1

@@ 1 Introduction
In addition to determining the mass and radius of PSR J0030+0451, Riley et al. (2019) produced, at the time, the most detailed hotspot map of the neutron star surface derived from observational data. The hotspots were found to have distinct shapes: one is approximately circular, while the other is like a crescent. Both of the hotspots were found to be located within the same hemisphere, suggesting magnetic structures beyond the regular dipole close to the star (Bilous et al., 2019). A few theoretical works followed, attempting to study the implication of the hotspot shapes on the magnetic field structure near the star (Chen et al., 2020; Kalapotharakos et al., 2021), and finding that a dipole plus quadrupole magnetic field can typically reproduce the observed X-ray light curve.
A recent investigation by Vinciguerra et al. (2024) revisited the hotspot modeling of PSR J0030+0451, finding the hotspots to be consistent with the polar caps of an off-center dipole magnetic field but with a nontrivial temperature distribution. In particular, the new study found that one of the hotspots has a high-temperature component near its edge. This study suggests that nontrivial temperature distribution across hotspots may need to be taken into account during the modeling process, which may significantly expand the parameter space. A theory that describes the temperature distribution across pulsar surface hotspots is urgently needed.
@@
In this paper, we attempt to take a first step towards a non-axisymmetric, physics-based model of the temperature distribution across hotspots on the surfaces of rotation-powered pulsars. Based on the theory of force-free pulsar magnetospheres developed by Gralla et al. (2017), we consider an off-center magnetic dipole and derive the distribution of the current density across the neutron star surface. Using this current distribution, we compute the surface heating rate using a simple prescription and derive the surface temperature distribution. Finally, we compute the pulsar’s light curve using the X-PSI global temperature module and incorporate it into pulse-profile modeling process. We perform a direct fit to the observed X-ray light-curve data of PSR J0030+0451 (J0030) and PSR J0437–4715 (J0437), aiming to showcase the explanatory power of our off-center dipole temperature map. This work illustrates the potential of adopting a physics-based hotspot configuration to model pulsar X-ray data, offering an alternative to the geometric methods currently in use (e.g., those proposed by Riley et al., 2019).
@@ 2 Method
Where $(J^{\hat{t}},J^{\hat{r}},J^{\hat{\theta}},J^{\hat{\phi}})$ are the components of the four-current components under the orthonormal basis $(\hat{\bm{t}},\hat{\bm{r}},\hat{\bm{\theta}},\hat{\bm{\phi}})$, and $J^{t}$ is given by:
@@
where $\Omega$ is the angular velocity, $\Omega_{Z}$ is the “frame-drag frequency” defined as $\Omega_{Z}=2\hat{I}\Omega/r^{3}$, and $\hat{I}$ is the moment of inertia, which defined as $\hat{I}=\mathcal{I}M_{\star}R_{\star}^{2}$, where we set the compactness $\mathcal{I}=2/5$. $M_{\star}$ and $R_{\star}$ are the mass and radius of the neutron star respectively. The function $\Lambda(\alpha,\beta)$ is a conserved quantity along each field line that determines the magnitude of the current density. Comparing with first-principles PIC simulations of the global pulsar magnetosphere, Gralla et al. (2017) were able to find an approximate expression for this function:
@@
This quantity $\alpha_{0}$ delineates the boundary of the polar cap, which is the region where open magnetic field lines originate from. The dipole magnetic moment $\mu$ is defined as $B_{0}\,R_{\star}^{3}$, where $B_{0}$ represents the magnetic field strength on the surface of the neutron star. In the subsequent sections, we adopt a value of $B_{0}=5\times 10^{8}\,\mathrm{G}$. Equation (9) is our definition of the pulsar polar cap for the rest of this paper, even when the designated regions are no longer geometrically similar to polar caps in the traditional sense.
@@
Given the four-current density across the polar cap, a heating prescription is required to compute the temperature distribution. This heating is intimately related to the dissipation of electromagnetic energy, particle acceleration, and the pair production process above the neutron star surface. As shown through analytic calculations by of Beloborodov (2008) and simulations by Timokhin & Arons (2013), particle acceleration at the polar cap depends on the ratio $j/c\rho_{\mathrm{GJ}}$, where $j$ is the magnitude of the three-current density and $\rho_{\mathrm{GJ}}$ is the Goldreich-Julian charge density Goldreich & Julian (1969). If this ratio is locally greater than 1 (the four-current is spacelike, $J^{2}>0$) or less than 0 ($j$ and $\rho_{\mathrm{GJ}}$ have opposite signs), these regions can build up a large voltage, accelerating particles to very high Lorentz factors, triggering $e^{\pm}$ pair production. A fraction of the produced pairs precipate on the polar cap, converting some of the energy dissipated in the gap to heat the neutron star surface. On the other hand, if $10<j/c\rho_{\mathrm{GJ}}<$ 1 (the four-current is timelike, $J^{2}<0$), the plasma can flow along the field line with a mildly relativistic Lorentz factor without building up a large voltage (Chen & Beloborodov, 2013).
@@
Figure 2: Pulse profile computed from the centered dipole field in three cases: Uniform hotspot temperature (red), Without return current (blue), With return current (green). All the other parameters to generate this plot are listed in Table 1. The magnetic inclination angle $\iota=45^{\circ}$, observer inclination angle $I=90^{\circ}$.
To compute the X-ray light curves from the surface temperature profiles, we use the open source X-PSI package by Riley et al. (2019) (https://github.com/xpsi-group/xpsi.git, Riley et al. (2023)), which follows a methodology developed over many previous works, including Bogdanov et al. (2019); Riley et al. (2019); Bogdanov et al. (2021); Riley et al. (2021). To describe the shape of the hotspots, the X-PSI package normally utilizes a model that represents the hot-spot surface through overlapping spherical caps with uniform temperature, but it also has the capability to accommodate irregular shapes of hot-spots by specifying the temperature distribution over the whole stellar surface  (see e.g. Das et al., 2024). It is the later mode that we are using in the calculations presented below. For definitiveness, the physical parameters we used for computing the light curves are summarized in Table 1.
@@ 3.1 Centered Dipole
Figure 3: 4-current density squared $J^{2}$ and temperature $T$ distribution of shifted hotspot projected on the star surface, with magnetic inclination $\iota=45^{\circ}$ and dipole shift $\delta=R_{\star}/2$ along different directions. Panels (a) and (b): $x$–direction shift; panels (c) and (d): $y$–direction; panels (e) and (f): $z$–direction shift. Each color map is normalized by the maximum value of $J^{2}$ and $T$.
Figure 4: Pulse profile computed from the shifted dipole field in three cases: $x$–direction shift (blue), $y$–direction shift (red) and $z$–direction shift (green), comparing with centered dipole field computation without including any shift (black dashed). All the other parameters to generate this plot are listed in Table 1 with magnetic inclination $\iota=45^{\circ}$, observer inclination $I=90^{\circ}$
@@ 3.2 Shifted Dipole
Figure 3 illustrates the effect of a shift $\delta=R_{\star}/2$ along the $x$, $y$, and $z$ directions on the polar cap 4-current density structure and temperature distribution. For example, in the panel (a), we present the $x$–direction shift field 4-current density distribution across the polar cap on the stellar surface. Compared to the centered dipole configuration, the north polar cap is greatly extended to a larger area, with a single hotspot at its center. In contrast, the south polar cap is compressed into a much smaller region. It should be noted that the small blank region at the center of the north pole is an artifact of the projection used in the plot. In reality, the hot spot does not exhibit any hollow feature. The panel (b) shows that the south polar cap , despite being much smaller, has a significantly higher temperature than the north. This is consistent with the expectation that regions located closer to the dipole center experience stronger magnetic fields and thus have stronger heating. Similar asymmetry between the temperatures of the two polar caps can be seen in the $z$–shifted dipole case as well (see panels (e) and (f) of Figure 3).
@@
The four-current density and temperature distribution for the $y$–shift scenario $\delta_{y}=R_{\star}/2$ in a dipole field are shown in the panels (c) and (d) of Figure 3. The two polar caps remain symmetric, and generally similar to the unshifted case shown in Figure 1. The main difference is that the hot spots become closer together compared to the centered dipole scenario. This displacement changes the distance between the two peaks in the X-ray light curve in one rotation period, as can be seen in Figure 4. Such a distinct outcome for a $y$–direction shift arises naturally because the star’s dipole moment lies in the $x-z$ plane, so shifting perpendicular to that plane yields behavior not seen in within-plane shifts.
@@
The panels (e) and (f) of Figure 3 show how a shift $\delta_{z}=R_{\star}/2$ along the $z$–direction affects the hotspot current density distribution and temperature map. In panel (e), the two polar caps display noticeably different areas, with the north polar cap undergoing an enlargement similar to what was observed in the $x$–shift scenario, and southern polar cap also exhibited a larger current density similar with what we see in $x$–direction shift. The panel (f) reveals that, in this configuration, the south hotspot appears significantly brighter, whereas the north hotspot—despite having a larger area— is comparatively fainter, similiar to the features seen in the $x$–shift case.
@@
The X-ray pulse profiles for shifts in the $x$–, $y$–, and $z$–directions are computed and presented in Figure 4 for a neutron star with magnetic inclination angle $\iota=45^{\circ}$, as seen by an observer at angle $I=90^{\circ}$ with respect to the rotation axis. Additional parameters, such as the mass and radius of the neutron star, are required to determine its external spacetime, thereby allowing us to generate the pulse profiles using the X-PSI package. We used the parameters of the pulsar J0030+0451 for these light curves, and the values are summarized in Table 1.
@@
Without any normalization added, our model indicates that for an $x$–direction shift, X-ray emission from the south polar cap dominates, yielding a significantly larger peak-to-valley ratio than the centered-dipole light curve (shown as the dashed curve). Although the south polar cap is smaller in area, it is brighter due to the locally enhanced magnetic field. Consequently, the maximum peak in the $x$–shift configuration is nearly twice that of the centered-dipole setup. In the $y$–direction shift scenario, the north and south polar caps move closer, merging into a single, highly boosted emission peak; Figure 4 shows that this configuration produces both the highest peak and the largest peak-to-valley ratio, implying a higher total flux than in the other shift cases. Finally, the $z$–direction shift presents an interesting contrast. Although the north polar cap appears fainter than the south cap in Figure 3, the observer inclination $I=90^{\circ}$ (equatorial viewing) brings the north cap closer to the line of sight. As a result, the pulse profile exhibits the smallest peak-to-valley ratio, and the south cap’s higher temperature does not substantially exceed the emission level seen in the centered-dipole configuration.
@@
Table 1: This is a summary for all the pulse-profile modeling parameters setting for (a) PSR J0030+0451 and (b) PSR J0437–4715. The J0030 parameters are also used in Section 3 to compute the hot spot temperature distribution of centered and shifted dipoles. Mass $M_{\star}$ Radius $R_{\star}$ refers the neutron star mass and radius, earth distance $D$ defined as the distance between this source and earth, $\nu$ is the Coordinate frequency of the mode of radiative asymmetry in the photosphere that is assumed to generate the pulsed signal [Hz]. $B$ in G [Gauss], is the surface magnetic field that take into the computation of magnetic moment.
@@
      observer inclination $I$       $\mathcal{U}(0,\frac{\pi}{2})$
@@
Table 2: Summary of Inference Model Parameters and Prior Settings: The observer inclination is denoted by $I$ and the magnetic inclination by $\iota$. The off-center dipole displacements along the $x$–, $y$–, and $z$–axes are represented by $x_{0}$, $y_{0}$, and $z_{0}$, respectively. $T_{s}$ denotes the temperature of the neutron star’s non-hotspot region (in arbitrary units). Here, $\mathcal{U}$ denotes a uniform distribution.
Figure 5: The posterior of the observer inclination $I$, magnetic inclination angle $\iota$ (both in radians), $x$$y$$z$ direction shifts and $\log_{10}T_{s}$ after applying constraints to the observed pulse profile of J0030. Blue is the posterior with Case 1: without return current. Red posterior is result in Case 2: with return current. The contour levels in the corner plot, going from deep to light colours, correspond to the 68%, 84% and 98.9% levels. The title of this plot indicates the median of the distribution as well as the range of the 68% credible interval. Here, $x$$y$$z$ shifts are measured in $\mathrm{km}$, while $T_{s}$ is in numerical units.
Highest Likelihood Case 1 Case 2
observer inclination $I$ 0.82 0.84
magnetic inclination $\iota$ 1.56 1.51
$x_{0}\,(\mathrm{km})$0.51 0.43
$y_{0}\,(\mathrm{km})$0.32 -0.35
$z_{0}\,(\mathrm{km})$ 0.31 0.66
$\log_{10}T_{s}$ 3.09 2.85

### v3

@@ 1 Introduction
In addition to determining the mass and radius of PSR J0030+0451, Riley et al. (2019); Miller et al. (2019) produced, at the time, the most detailed hotspot map of the neutron star surface derived from observational data. The hotspots were found to have distinct shapes: one is approximately circular, while the other is like a crescent. Both of the hotspots were found to be located within the same hemisphere, suggesting magnetic structures beyond the regular dipole close to the star (Bilous et al., 2019). A few theoretical works followed, attempting to study the implication of the hotspot shapes on the magnetic field structure near the star (Chen et al., 2020; Kalapotharakos et al., 2021), and finding that a dipole plus quadrupole magnetic field can typically reproduce the observed X-ray light curve.
A recent investigation by Vinciguerra et al. (2024) revisited the hotspot modeling of PSR J0030+0451, finding the hotspots to be consistent with the polar caps of an off-center dipole magnetic field but with a nontrivial temperature distribution due to the implementation of new background constraints. In particular, the new study found that one of the hotspots has a high-temperature component near its edge. This study suggests that nontrivial temperature distribution across hotspots may need to be taken into account during the modeling process, which may significantly expand the parameter space. A theory that describes the temperature distribution across pulsar surface hotspots is urgently needed.
@@
In this paper, we attempt to take a first step towards a non-axisymmetric, physics-based model of the temperature distribution across hotspots on the surfaces of rotation-powered pulsars. Based on the theory of force-free pulsar magnetospheres developed by Gralla et al. (2017), we consider an off-center magnetic dipole and derive the distribution of the current density across the neutron star surface. Using this current distribution, we compute the surface heating rate using a simple prescription and derive the surface temperature distribution. Finally, we compute the pulsar’s light curve using the X-PSI global temperature module and incorporate it into pulse-profile modeling process. We perform a direct fit to the observed X-ray light-curve data of PSR J0030+0451 (J0030) (Bogdanov et al., 2019) and PSR J0437–4715 (J0437) (Choudhury et al., 2024), aiming to showcase the explanatory power of our off-center dipole temperature map. This work illustrates the potential of adopting a physics-based hotspot configuration to model pulsar X-ray data, offering an alternative to the geometric methods currently in use (e.g., those proposed by Riley et al., 2019).
@@ 2 Method
[…]
@@
where $\Omega$ is the angular velocity, $\Omega_{Z}$ is the “frame-drag frequency” defined as $\Omega_{Z}=2\hat{I}\Omega/r^{3}$, and $\hat{I}$ is the moment of inertia, which defined as $\hat{I}=\mathcal{I}M_{\star}R_{\star}^{2}$, where we set the compactness $\mathcal{I}=2/5$. $M_{\star}$ and $R_{\star}$ are the mass and radius of the neutron star respectively. The normalized time component $J^{\hat{t}}$ is the force-free charge density, and can be identified with the Goldreich-Julian charge density $\rho_{\mathrm{GJ}}$ first defined by Goldreich & Julian (1969). The function $\Lambda(\alpha,\beta)$ is a conserved quantity along each field line that determines the magnitude of the current density. Comparing with first-principles PIC simulations of the global pulsar magnetosphere, Gralla et al. (2017) were able to find an approximate expression for this function:
@@
This quantity $\alpha_{0}$ delineates the boundary of the polar cap, which is the region where open magnetic field lines originate from. The dipole magnetic moment $\mu$ is defined as $B_{0}\,R_{\star}^{3}$, where $B_{0}$ represents the magnetic field strength on the surface of the neutron star. In the subsequent sections, we adopt a value of $B_{0}=5\times 10^{8}\,\mathrm{G}$ based on the rough spin-down estimation for millisecond pulsar populations. Equation (9) is our definition of the pulsar polar cap for the rest of this paper, even when the designated regions are no longer geometrically similar to polar caps in the traditional sense.
@@
Given the four-current density across the polar cap, a heating prescription is required to compute the temperature distribution. This heating is intimately related to the dissipation of electromagnetic energy, particle acceleration, and the pair production process above the neutron star surface. As shown through analytic calculations by of Beloborodov (2008) and simulations by Timokhin & Arons (2013), particle acceleration at the polar cap depends on the ratio $j/c\rho_{\mathrm{GJ}}$, where $j$ is the magnitude of the three-current density defined in Equations (4)–(6), and $\rho_{\mathrm{GJ}}$ is the Goldreich-Julian charge density, which is equivalent to $J^{\hat{t}}$ defined in Equation (3). If this ratio is locally greater than 1 (the four-current is spacelike, $J^{2}>0$) or less than 0 ($j$ and $\rho_{\mathrm{GJ}}$ have opposite signs), these regions can build up a large voltage, accelerating particles to very high Lorentz factors, triggering $e^{\pm}$ pair production. A fraction of the produced pairs precipate on the polar cap, converting some of the energy dissipated in the gap to heat the neutron star surface. On the other hand, if $0<j/c\rho_{\mathrm{GJ}}<1$ (the four-current is timelike, $J^{2}<0$), the plasma can flow along the field line with a mildly relativistic Lorentz factor without building up a large voltage (Chen & Beloborodov, 2013).
@@
Figure 2: Pulse profile of bolometric flux computed from the centered dipole field in three cases: Uniform hotspot temperature (red), Without return current (blue), With return current (green). All the other parameters to generate this plot are listed in Table 1. The magnetic inclination angle $\iota=45^{\circ}$, observer inclination angle $I=90^{\circ}$.
To compute the X-ray light curves from the surface temperature profiles, we use the open source X-PSI package used by  Riley et al. (2019) (https://github.com/xpsi-group/xpsi.git, Riley et al. (2023)), which follows a methodology developed over many previous works (Riley, 2019). This code was bechmarked in Bogdanov et al. (2019, 2021), and used in analysis Riley et al. (2019, 2021). To describe the shape of the hotspots, the X-PSI package normally utilizes a model that represents the hot-spot surface through overlapping spherical caps with uniform temperature, but it also has the capability to accommodate irregular shapes of hot-spots by specifying the temperature distribution over the whole stellar surface  (see e.g. Das et al., 2024). It is the later mode that we are using in the calculations presented below. For definitiveness, the physical parameters we used for computing the light curves are summarized in Table 1.
@@ 3.1 Centered Dipole
Figure 3: 4-current density squared $J^{2}$ and temperature $T$ distribution of shifted hotspot projected on the star surface, with magnetic inclination $\iota=45^{\circ}$ and dipole shift $\delta=R_{\star}/2$ along different directions. Panels (a) and (b): xx–direction shift; panels (c) and (d): yy–direction; panels (e) and (f): zz–direction shift. Each color map is normalized by the maximum value of $J^{2}$ and $T$.
Figure 4: Pulse profile computed from the shifted dipole field in three cases: xx–direction shift (blue), yy–direction shift (red) and zz–direction shift (green), comparing with centered dipole field computation without including any shift (black dashed). All the other parameters to generate this plot are listed in Table 1 with magnetic inclination $\iota=45^{\circ}$, observer inclination $I=90^{\circ}$
@@ 3.2 Shifted Dipole
Figure 3 illustrates the effect of a shift $\delta=R_{\star}/2$ along the $x$, $y$, and $z$ directions on the polar cap 4-current density structure and temperature distribution. For example, in the panel (a), we present the xx–direction shift field 4-current density distribution across the polar cap on the stellar surface. Compared to the centered dipole configuration, the north polar cap is greatly extended to a larger area, with a single hotspot at its center. In contrast, the south polar cap is compressed into a much smaller region. It should be noted that the small blank region at the center of the north pole is an artifact of the projection used in the plot. In reality, the hot spot does not exhibit any hollow feature. The panel (b) shows that the south polar cap , despite being much smaller, has a significantly higher temperature than the north. This follows from the fact that, in our model, the local surface temperature is set by the magnitude of the current density $j$. Each open magnetic flux surface carries a conserved current $\Lambda(\alpha,\beta)$, fixed in the far zone where the field is symmetric across the current sheet. Because both caps intercept the same total current, squeezing one cap into a smaller area inevitably raises $j$ there, and thus (after the red-shift factor $\sqrt{1-2M/R_{\star}}$ is applied) delivers more heating per unit area. Similar asymmetry between the temperatures of the two polar caps can be seen in the zz–shifted dipole case as well (see panels (e) and (f) of Figure 3).
@@
The four-current density and temperature distribution for the yy–shift scenario $\delta_{y}=R_{\star}/2$ in a dipole field are shown in the panels (c) and (d) of Figure 3. The two polar caps remain symmetric, and generally similar to the unshifted case shown in Figure 1. The main difference is that the hot spots become closer together compared to the centered dipole scenario. This displacement changes the distance between the two peaks in the X-ray light curve in one rotation period, as can be seen in Figure 4. Such a distinct outcome for a yy–direction shift arises naturally because the star’s dipole moment lies in the xzx-z plane, so shifting perpendicular to that plane yields behavior not seen in within-plane shifts.
@@
The panels (e) and (f) of Figure 3 show how a shift $\delta_{z}=R_{\star}/2$ along the zz–direction affects the hotspot current density distribution and temperature map. In panel (e), the two polar caps display noticeably different areas, with the north polar cap undergoing an enlargement similar to what was observed in the xx–shift scenario, and southern polar cap also exhibited a larger current density similar with what we see in xx–direction shift. The panel (f) reveals that, in this configuration, the south hotspot appears significantly brighter, whereas the north hotspot—despite having a larger area— is comparatively fainter, similiar to the features seen in the xx–shift case.
@@
The X-ray pulse profiles for shifts in the xx–, yy–, and zz–directions are computed and presented in Figure 4 for a neutron star with magnetic inclination angle $\iota=45^{\circ}$, as seen by an observer at angle $I=90^{\circ}$ with respect to the rotation axis. Additional parameters, such as the mass and radius of the neutron star, are required to determine its external spacetime, thereby allowing us to generate the pulse profiles using the X-PSI package. We used the parameters of the pulsar J0030+0451 for these light curves, and the values are summarized in Table 1.
@@
Without any normalization added, our model indicates that for an xx–direction shift, X-ray emission from the south polar cap dominates, yielding a significantly larger peak-to-valley ratio than the centered-dipole light curve (shown as the dashed curve). Although the south polar cap is smaller in area, it is brighter due to the locally enhanced magnetic field. Consequently, the maximum peak in the xx–shift configuration is nearly twice that of the centered-dipole setup. In the yy–direction shift scenario, the north and south polar caps move closer, merging into a single, highly boosted emission peak; Figure 4 shows that this configuration produces both the highest peak and the largest peak-to-valley ratio, implying a higher total flux than in the other shift cases. Finally, the zz–direction shift presents an interesting contrast. Although the north polar cap appears fainter than the south cap in Figure 3, the observer inclination $I=90^{\circ}$ (equatorial viewing) brings the north cap closer to the line of sight. As a result, the pulse profile exhibits the smallest peak-to-valley ratio, and the south cap’s higher temperature does not substantially exceed the emission level seen in the centered-dipole configuration.
@@
Table 1: This is a summary for all the pulse-profile modeling parameters setting for (a) PSR J0030+0451 and (b) PSR J0437–4715. The J0030 parameters are also used in Section 3 to compute the hot spot temperature distribution of centered and shifted dipoles. Mass $M_{\star}$ Radius $R_{\star}$ refers the neutron star mass and radius, earth distance $D$ defined as the distance between this source and earth, $\nu$ is the Coordinate frequency of the mode of radiative asymmetry in the photosphere that is assumed to generate the pulsed signal [Hz]. $B$ in G [Gauss], is the surface magnetic field that take into the computation of magnetic moment. The frequency and earth distance of J0030 are derived from radio observation Arzoumanian et al. (2018), and the frequency and earth distance of J0437 was from Reardon et al. (2024).
@@
  observer inclination II   $\mathcal{U}(0,\frac{\pi}{2})$
@@
Table 2: Summary of Inference Model Parameters and Prior Settings: The observer inclination is denoted by II and the magnetic inclination by $\iota$. The off-center dipole displacements along the xx–, yy–, and zz–axes are represented by $x_{0}$, $y_{0}$, and $z_{0}$, respectively. $T_{s}$ denotes the temperature of the neutron star’s non-hotspot region (in arbitrary units). Here, $\mathcal{U}$ denotes a uniform distribution.
Figure 5: The posterior of the observer inclination $I$, magnetic inclination angle $\iota$ (both in radians), xxyyzz direction shifts and $\log_{10}T_{s}$ after applying constraints to the observed pulse profile of J0030. Blue is the posterior with Case 1: without return current. Red posterior is result in Case 2: with return current. The contour levels in the corner plot, going from deep to light colours, correspond to the 68%, 84% and 98.9% levels. The title of this plot indicates the median of the distribution as well as the range of the 68% credible interval. Here, xxyyzz shifts are measured in $\mathrm{km}$, while $T_{s}$ is in numerical units.
Maximum Likelihood Case 1 Case 2
observer inclination II 0.78 0.68
magnetic inclination $\iota$ 1.52 0.98
$x_{0}\,(\mathrm{km})$ -0.35 0.66
$y_{0}\,(\mathrm{km})$ -0.22 -0.29
$z_{0}\,(\mathrm{km})$ -0.25 -0.16
$\log_{10}T_{s}$ 2.40 2.62